OpenStudy (anonymous):

how to solve x^2=-27

OpenStudy (anonymous):

Is that the whole question?

OpenStudy (anonymous):

I believe that this is a quadratic equation... Just a minute..

OpenStudy (anonymous):

To solve the quadratic equation: x^2 = 27 We'll take square root both sides: sqrt x^2 = sqrt 27 We'll write 27 = 3^3 = 3*3^2 |x| = 3sqrt3 Does this help?

OpenStudy (anonymous):

so the last answer will be the correct answser

OpenStudy (anonymous):

Just take the square root of both sides. So you have \[x=\pm \sqrt{-27}\]Since you have a negative 27. You have to use complex numbers to solve. Note that\[i=\sqrt{-1}\] So you have \[x=\pm(\sqrt{-1}*\sqrt{3}*\sqrt{9})\] This reduces to\[x=\pm3i \sqrt{3}\]